40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15324     Entry time: Mon May 11 00:12:34 2020
Author: gautam 
Type: Update 
Category: LSC 
Subject: RF only PRFPMI 

Finally - Attachment #1. This plot uses 16 Hz EPICS data. All y-axes are uncalibrated for now, but TRX/TRY are normalized such that the POX/POY lock yields a transmission of 1. CARM UGF is only ~3 kHz, no boosts were turned on yet. 

Attachment #2 and Attachment #3 are phone photos of the camera images of the various ports. After some alignment work, the transmitted arm powers were ~200, i.e. PRG ~10. fwiw, this is the darkest i've ever seen the 40m dark port. c.f. 2016. Of course, the exposure time / ND filter / light levels could all have changed.

This work was possible during the daytime (~6pm PDT), but probably only because it was Sunday. The other rate limiting factor here is the franky terrible IMC duty cycle. TBH, I didn't honestly expect to get so far and ran out of time, but I think the next steps are:

  1. Turn on some sensing lines and calibrate CARM/DARM.
  2. Transition vertex control to 1f signals.
  3. Whiten DARM.
  4. Turn on some ASC for better power stabilization.
  5. Scan the CARM offset and check that we are truly on resonance
  6. Noise budget.

As usual, I would like to request that we don't change the IFO as far as possible until the BHD vent, i found it pretty difficult to get here.


Attachment #4 now shows the measured DARM OLTF when DARM is entirely on AS55_Q control. UGF is ~120 Hz and the phase margin is ~30 deg, seems okay for a first attempt. I'll now need to infer the OLTF over a wider range of frequencies by lining this measurement up with some model, so that I can undo the loop in plotting the DARM ASD.

Attachment 1: PRFPMIlock.pdf  62 kB  Uploaded Mon May 11 01:35:30 2020  | Hide | Hide all
PRFPMIlock.pdf
Attachment 2: IMG_8549.JPG  1.095 MB  Uploaded Mon May 11 01:35:42 2020  | Hide | Hide all
IMG_8549.JPG
Attachment 3: IMG_8548.JPG  881 kB  Uploaded Mon May 11 01:35:49 2020  | Hide | Hide all
IMG_8548.JPG
Attachment 4: DARM_OLTF.pdf  86 kB  Uploaded Wed May 13 00:00:11 2020  | Hide | Hide all
DARM_OLTF.pdf
ELOG V3.1.3-